OpenStudy (anonymous):

Graph the function. Label the vertex and the axis of symmetry. g(x)=3/4x^2+3x+9/4

OpenStudy (tamtoan):

again, set g(x) = 0, multiply every thing by 4 to get rid of the 4 in the bottom, now find vertex like the other problem...from that you can get a few more points on both side and graph

OpenStudy (anonymous):

so the new eqtn is 3x^2+12x+36 ?

OpenStudy (tamtoan):

9/4 multiply 4 you will get 9 :)...from there you got your a, b and c

OpenStudy (anonymous):

oh yeah lol i forgot it cancels the denominator out haha . so its 3x^2+12x+9 ?

OpenStudy (tamtoan):

yeah, also if 3x^2 +12x + 9 = 0...you can see that there's a 3 factor in all terms you can divide to make the number smaller, hence, make it easier in calculation

OpenStudy (anonymous):

yea i just divided all them by 3. so to find points for the graph can i just pick any number? like say i pick 0 . x^2+4x+3=y . sooo , x^2+4(0)+3=y 3=y so a points can be (0,3) ?

OpenStudy (anonymous):

0^2**

OpenStudy (tamtoan):

you know this parabola point upward right..now what is your vertex ?

OpenStudy (anonymous):

yea i know that , vertex is (-2,0) right?

OpenStudy (tamtoan):

if you sub x = -2 in, do you get y = 0 ? or something else?

OpenStudy (anonymous):

-1.

OpenStudy (anonymous):

oh ok i think i get this

OpenStudy (anonymous):

so the vertex is (-2,-1) . right . & i found the points (-1,0) and point (0,3) .

OpenStudy (tamtoan):

now that is your vertex, and x = -2 is the axis of symetry..you let x = -1 find y, let x = -3 find y...then you have 3 points you can graph, if you want to be more accurate, you can get 2 more values , one on each side of the axis of symetry, and then...done